OpenStudy (fatimab2000):

PLEASE HELP Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6) A) 7 B) -14 C) 0 D) 14

OpenStudy (sweetburger):

y2-y1/x2-x1 = m

OpenStudy (dayakar):

d is the right answer

OpenStudy (dayakar):

slope(m) = 8-(-6)/-3+4 = 14

OpenStudy (texaschic101):

you will use the slope formula : slope = (y2 - y1) / (x2 - x1) and you have 2 sets of points.. set 1 : (-4,8)....so we will call these points : x1 = -4 and y1 = 8 set 2 : (-3,-6)...and we will call these points : x2 = -3 and y2 = -6 now we sub into the slope formula...pay close attention to your signs.. slope =(-6 - 8) / (-3 - (-4) slope = -14 / (-3 + 4) slope = -14/1....which is the same as -14